# presep= ]e1,e2,e3

### v1

@@ 6.1 Description of dataset
The West Texas Intermediate (WTI) is a grade of crude oil and is considered one of the main benchmarks in oil pricing. We are interested to examine whether the spot price of WTI is conditionally regularly varying given the Crude Oil Volatility Index (CBOE OVX) and, if this is the case, how the tail index is influenced by different levels of implied volatility. The CBOE OVX is constructed by applying the VIX (Chicago Board Options Exchange’s Volatility Index) methodology (see Whaley (2000)) to options on the United States Oil Fund (USO), using a weighted average of mid-quote prices from calls and puts across a range of strike prices. This aggregated implied volatility is extrapolated to a 30‐day horizon to provide a forward-looking measure of the expectation of the market of price fluctuations and is expressed in percent.
Now, as USO is designed to track the WTI price using future contracts, the OVX provides a forward looking market expectation of the WTI price volatility. Thus, in the case of regularly varying marginal distributions of the WTI, we expect that the tail index varies as a function of OVX. While the OVX is mean-reverting and can be considered stationary (see Chen, Liu and Li (2018)), oil prices are generally not stationary, and are consequently often transformed to log-returns (see Hamilton (1983); Sadorsky (1999)) to achieve stationarity. We follow this approach. The corresponding time series are depicted in Figure 4. Note how high activity periods such as the financial crisis (2008), oil price collapse due to US and OPEC oil war (2015), covid period (2020-2021), and the Russian and Ukrainian conflict (2022), are visually observable. Remark that the WTI price attained a negative value on the 20th of April 2020, however this observation has been removed in order to transform with logarithms.
@@ 6.2 Analysis of crude oil spot price
Figure 6: Plots of $\widehat{\gamma}_{n}(x)$, using the negative (left) and positive (right) WTI log-returns as $(Y_{n})$ and the uniformly transformed OVX as $(X_{n})$ (covariate process). In the top panels we depict the conditional Hill estimator, as of function of $k_{n}$, i.e. the number of top order statistics included, for four chosen covariate values $x\in\{0.201,0.422,0.582,0.995\}$ (dotted, dashed, long-dashed and solid, respectively). In the bottom panels $\widehat{\gamma}_{n}(x)$ is plotted as a function of $x$ for four fixed number of order statistics, $k_{n}\in\{80,241,552,959\}$ (dotted, dashed, long-dashed and solid, respectively).
In Figure 6 we consider the conditional Hill estimator of the WTI log-returns given the uniformly transformed OVX. We observe that the largest $\widehat{\gamma}_{n}(x)$ value is attained for both negative and positive sides of the WTI log-returns, when the implied volatility is large. This is in accordance with the common financial intuition that large implied volatility implies heavier tails (and thus fewer moments) for log-returns of markets on which it is based. Furthermore, when the implied volatility is large, the negative side of the WTI log-returns exhibits slightly heavier tails than the positive side. This has been documented for many assets (see Christie (1982)), where shock effects can create large upswings but even greater downswings. Since the conditional Hill estimator can be viewed as a function of the covariate value $x$ for a fixed sample fraction $k_{n}$ bivariately, we also plot these “risk profiles”, where slightly noisy smiles can be seen, which indicates that the tail index is largest when volatility is low or high.
@@ 7 Conclusion and outlook
Apart from considering stronger intertemporal dependence structures, our work also allows us to consider extensions in several directions. First, having an estimate of the entire tail at our disposal our results can be used to construct and statistically analyse conditional risk measures such as conditional expected shortfall or conditional tail moments. Secondly, a promising direction is to consider specific models where the same time series can be used as the covariate and response, but at different lags; Markov processes are particularly attractive in this connection. Finally, removing top observations can sometimes lead to the construction of stable estimators that allow transparent threshold selection as in Bladt, Albrecher and Beirlant (2020), which is yet to be explored both for univariate and conditional cases.
@@ A.1 Proof of Theorem 1 (Tail consistency, $m$-dependence)
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]\to s^{-\alpha_{x}}g(x),\ s>0.$
@@
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]=g(x)\frac{\overline{F}^{x}(su_{n})}{\overline{F}^{x}(u_{n})}\int(su_{n})^{-(\alpha_{x-uh_{n}}-\alpha_{x})}\frac{L^{x-uh_{n}}(su_{n})}{L^{x}(su_{n})}\frac{g(x-h_{n}u)}{g(x)}K\left(u\right)\mathrm{d}u.$
@@
$(su_{n})^{-(\alpha_{x-uh_{n}}-\alpha_{x})}=\exp\left(-(\alpha_{x-uh_{n}}-\alpha_{x})\log(su_{n})\right)$
$\leq\exp\left(c_{\alpha}h_{n}\lvert u\rvert\log(su_{n})\right)=\exp\left(c_{\alpha}\lvert u\rvert h_{n}\log(u_{n})\right)\exp\left(c_{\alpha}h_{n}\lvert u\rvert\log(s)\right)\to 1,$
@@
$(su_{n})^{-(\alpha_{x-uh_{n}}-\alpha_{x})}\geq\exp\left(-c_{\alpha}\lvert u\rvert h_{n}\log(u_{n})\right)\exp\left(-c_{\alpha}h_{n}\lvert u\rvert\log(s)\right)\to 1,$
uniformly in $u\in[-1,1]$, using condition 1 and $h_{n}\log(u_{n})\to 0$. Now, condition 3 implies
$\exp\big{(}-\log(su_{n})h_{n}c_{L}\left|u\right|\big{)}\leq\frac{L^{x-uh_{n}}(su_{n})}{L^{x}(su_{n})}\leq\exp\big{(}\log(su_{n})h_{n}c_{L}\left|u\right|\big{)}$
@@
By the dominated convergence theorem get get $\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]\to g(x)s^{-\alpha_{x}}$. If $g(x)=0$, similar arguments, using Lipschitz continuity of $g$, yield the result. ∎
@@
$\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}t\right\}}K^{q}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}\to t^{-\alpha_{x}}g(x)\int K^{q}\left(u\right)\mathrm{d}u.$
@@
$\widetilde{T}^{x}_{n}(s)=\widetilde{V}^{x}_{n}(s)/g_{n}(x)$, it remains to show that $\widetilde{V}^{x}_{n}(s)\overset{\mathrm{Pr}}{\to}g(x)s^{-\alpha_{x}}$.
@@
$\widetilde{V}^{x,i}_{n}(s)$ $=\frac{m}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{j=1}^{\left\lfloor\frac{n}{m}\right\rfloor}K\left(\frac{x-X_{(j-1)m+i}}{h_{n}}\right)1_{\left\{Y_{(j-1)m+i}>su_{n}\right\}},$
$R_{n}(s)$ $=\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}.$
@@
$\mathbb{E}\left[\widetilde{V}^{x,i}_{n}(s)\right]=\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}s\right\}}K\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}\to s^{-\alpha_{x}}g(x),$
@@
$\operatorname{var}\left(\widetilde{V}^{x,i}_{n}(s)\right)=\frac{\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor m}{nh_{n}\overline{F}^{x}(u_{n})}\Bigg{[}\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}s\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}-\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}s\right\}}K\left(\frac{x-X}{h_{n}}\right)\right]^{2}}{h_{n}\overline{F}^{x}(u_{n})}\Bigg{]},$
@@
$\mathbb{E}\left[R_{n}(s)\right]=\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\frac{1}{n}\frac{\mathbb{E}\left[K\left(\frac{x-X}{h_{n}}\right)1_{\left\{Y>su_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}\to 0,$
$\mathbb{E}\left[R_{n}(s)^{2}\right]=\frac{1}{\left(nh_{n}\overline{F}^{x}(u_{n})\right)^{2}}\biggl{(}\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]$
$\qquad\qquad\quad\ \ +2\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}\sum_{i=j}^{n}\mathbb{E}\left[1_{\left\{Y_{i}>u_{n}s\right\}}K\left(\frac{x-X_{i}}{h_{n}}\right)1_{\left\{Y_{j}>u_{n}s\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)\right]\biggr{)}$
$\leq\frac{\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)}{\left(nh_{n}\overline{F}^{x}(u_{n})\right)^{2}}\biggl{(}\mathbb{E}\left[K^{2}\left(\frac{x-X}{h_{n}}\right)1_{\left\{Y>su_{n}\right\}}\right]+2\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\sqrt{\mathbb{E}\left[1_{\left\{Y>u_{n}s\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]^{2}}\ \biggr{)}$
$=\frac{\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)}{n^{2}h_{n}\overline{F}^{x}(u_{n})}\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}t\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}\left(1+2\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\right)\to 0,$
@@
A.2 Proof of Theorem 2 (Tail consistency, $\alpha$-mixing)
@@
$\text{var}(\widetilde{V}_{n}^{x}(s))$ $=\frac{1}{(nh_{n}\overline{F}^{x}(u_{n}))^{2}}\Bigg{(}\sum_{j=1}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]-n^{2}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]^{2}$
$\qquad\qquad\qquad+2n\sum_{j=1}^{n}\left(1-\frac{j}{n}\right)\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]\Bigg{)}$
$\leq\frac{1}{(nh_{n}\overline{F}^{x}(u_{n}))^{2}}\Bigg{(}\sum_{j=1}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]$
$\quad+2n\sum_{j=1}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right)\Bigg{)},$
@@
$\frac{1}{h_{n}\overline{F}^{x}(u_{n})}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]$
$=h_{n}\int_{[-1,1]^{2}}\biggl{\{}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}\mid X_{0}=x+w_{1}h_{n},X_{j}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n})}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr{\}}\mathrm{d}(w_{1},w_{2})=\mathcal{O}(h_{n}),$
where the integral is bounded, for $n$ large enough, due to condition 6 and smoothness of $g$. Now, Billingsley’s inequality with the $\alpha$-mixing assumption yields that
$\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right)\right|$
$\leq 4\alpha(j)\left\lVert K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right\rVert_{\infty}\left\lVert K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right\rVert_{\infty}=\mathcal{O}(\alpha(j)),$
using that the kernel is bounded. Let $d_{n}=\lceil h_{n}^{-c_{x}}\rceil$ for some $c_{x}\geq 0$. Combing the above considerations yields that
$\frac{2n}{(nh_{n}\overline{F}^{x}(u_{n}))^{2}}\left|\sum_{j=1}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right)\right|$
$\leq\frac{\text{cst}}{n(h_{n}\overline{F}^{x}(u_{n}))^{2}}\left(\sum_{j=1}^{d_{n}-1}h_{n}^{2}\overline{F}^{x}(u_{n})+\sum_{j=d_{n}}^{\infty}j^{-\eta}\right)\leq\frac{\text{cst}}{n(h_{n}\overline{F}^{x}(u_{n}))^{2}}\left(d_{n}h_{n}^{2}\overline{F}^{x}(u_{n})+d_{n}^{-\eta+1}\right)$
$=\text{cst}\left(\frac{d_{n}}{n\overline{F}^{x}(u_{n})}+\frac{d_{n}^{-\eta+1}}{n(h_{n}\overline{F}^{x}(u_{n}))^{2}}\right)\sim\text{cst}\left(\frac{1}{nh_{n}^{c_{x}}\overline{F}^{x}(u_{n})}+\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}))^{2}}\right)\to 0.$

### v2

@@ 6.1 Description of dataset
The West Texas Intermediate (WTI) is a grade of crude oil and is considered one of the main benchmarks in oil pricing. We are interested to examine whether the spot price of WTI is conditionally regularly varying given the Crude Oil Volatility Index (CBOE OVX) and, if this is the case, how the tail index is influenced by different levels of implied volatility. The CBOE OVX is constructed by applying the VIX (Chicago Board Options Exchange’s Volatility Index) methodology (see whaley2000investor) to options on the United States Oil Fund (USO), using a weighted average of mid-quote prices from calls and puts across a range of strike prices. This aggregated implied volatility is extrapolated to a 30‐day horizon to provide a forward-looking measure of the expectation of the market of price fluctuations and is expressed in percent.
Now, as USO is designed to track the WTI price using future contracts, the OVX provides a forward looking market expectation of the WTI price volatility. Thus, in the case of regularly varying marginal distributions of the WTI, we expect that the tail index varies as a function of OVX. While the OVX is mean-reverting and can be considered stationary (see chen2018predictive), oil prices are generally not stationary, and are consequently often transformed to log-returns (see hamilton1983oil; sadorsky1999oil) to achieve stationarity. We follow this approach. The corresponding time series are depicted in Figure 4. Note how high activity periods such as the financial crisis (2008), oil price collapse due to US and OPEC oil war (2015), covid period (2020-2021), and the Russian and Ukrainian conflict (2022), are visually observable. Remark that the WTI price attained a negative value on the 20th of April 2020, however this observation has been removed in order to transform with logarithms.
@@ 6.2 Analysis of crude oil spot price
Figure 6: Plots of $\widehat{\gamma}_{n}(x)$, using the negative (left) and positive (right) WTI log-returns as $(Y_{n})$ and the uniformly transformed OVX as $(X_{n})$ (covariate process). In the top panels we depict the conditional Hill estimator, as of function of $k_{n}$, i.e. the number of top order statistics included, for four chosen covariate values $x\in\{0.231,0.402,0.995\}$ (red, green, and blue, respectively). In the bottom panels $\widehat{\gamma}_{n}(x)$ is plotted as a function of $x$ for four fixed number of order statistics, $k_{n}\in\{80,241,552,959\}$ (dotted, dashed, long-dashed and solid, respectively).
In Figure 6 we consider the conditional Hill estimator of the WTI log-returns given the uniformly transformed OVX. We observe that the largest $\widehat{\gamma}_{n}(x)$ value is attained for both negative and positive sides of the WTI log-returns, when the implied volatility is large. This is in accordance with the common financial intuition that large implied volatility implies heavier tails (and thus fewer moments) for log-returns of markets on which it is based. Furthermore, when the implied volatility is large, the negative side of the WTI log-returns exhibits slightly heavier tails than the positive side. This has been documented for many assets (see Christie_82), where shock effects can create large upswings but even greater downswings. In addition, the confidence intervals that arise from Theorem 5 are depicted. For the both sides, the lower bound of the estimator corresponding to high volatility is separated from the upper bound of the of the estimators, corresponding to low and mid volatility, within the stable regions. This supports the claim the the the WTI log-returns change tail index according the implied volatility. Since the conditional Hill estimator can be viewed as a function of the covariate value $x$ for a fixed sample fraction $k_{n}$ bivariately, we also plot these “risk profiles”, where slightly noisy right-skewed smiles can be seen, which indicates that the tail index is large when volatility is low and even larger when high.
@@ 7 Conclusion and outlook
Apart from considering stronger intertemporal dependence structures, our work also allows us to consider extensions in several directions. First, having an estimate of the entire tail at our disposal our results can be used to construct and statistically analyse conditional risk measures such as conditional expected shortfall or conditional tail moments. Secondly, a promising direction is to consider specific models where the same time series can be used as the covariate and response, but at different lags; Markov processes are particularly attractive in this connection. Finally, removing top observations can sometimes lead to the construction of stable estimators that allow transparent threshold selection as in bladt2020threshold, which is yet to be explored both for univariate and conditional cases.
@@
Data availability statement
The data were accessed through a subscription-based platform and are not publicly available.
@@ Appendix A Proofs of main results from Section 3
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]\to s^{-1/\gamma(x)}g(x),\ s>0.$
@@
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]=g(x)\frac{\overline{F}^{x}(su_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\int(su_{n}^{x})^{-(1/\gamma(x-uh_{n})-1/\gamma(x))}\frac{L^{x-uh_{n}}(su_{n}^{x})}{L^{x}(su_{n}^{x})}\frac{g(x-h_{n}u)}{g(x)}K\left(u\right)\mathrm{d}u.$
@@
$(su_{n}^{x})^{-(1/\gamma(x-uh_{n})-1/\gamma(x))}=\exp\left(-(1/\gamma(x-uh_{n})-1/\gamma(x))\log(su_{n}^{x})\right)$
$\leq\exp\left(c_{\gamma}h_{n}\lvert u\rvert\log(su_{n}^{x})\right)=\exp\left(c_{\gamma}\lvert u\rvert h_{n}\log(u_{n}^{x})\right)\exp\left(c_{\gamma}h_{n}\lvert u\rvert\log(s)\right)\to 1,$
@@
$(su_{n}^{x})^{-(1/\gamma(x-uh_{n})-1/\gamma(x))}\geq\exp\left(-c_{\gamma}\lvert u\rvert h_{n}\log(u_{n}^{x})\right)\exp\left(-c_{\gamma}h_{n}\lvert u\rvert\log(s)\right)\to 1,$
uniformly in $u\in[-1,1]$, using condition 1 and $h_{n}\log(u_{n}^{x})\to 0$. Now, condition 3 implies
$\exp\big(-\log(su_{n}^{x})h_{n}c_{L}\left|u\right|\big)\leq\frac{L^{x-uh_{n}}(su_{n}^{x})}{L^{x}(su_{n}^{x})}\leq\exp\big(\log(su_{n}^{x})h_{n}c_{L}\left|u\right|\big)$
@@
By the dominated convergence theorem get $\mathbb{E}[\widetilde{V}^{x}_{n}(s)]\to g(x)s^{-1/\gamma(x)}$. If $g(x)=0$, similar arguments, using Lipschitz continuity of $g$, yield the result. ∎
@@
$\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}t\right\}}K^{q}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\to t^{-1/\gamma(x)}g(x)\int K^{q}\left(u\right)\mathrm{d}u.$
@@
$\widetilde{T}^{x}_{n}(s)=\widetilde{V}^{x}_{n}(s)/g_{n}(x)$, it remains to show that $\widetilde{V}^{x}_{n}(s)\overset{\mathrm{Pr}}{\to}g(x)s^{-1/\gamma(x)}$.
@@
$\widetilde{V}^{x,i}_{n}(s)$ $=\frac{m}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=1}^{\left\lfloor\frac{n}{m}\right\rfloor}K\left(\frac{x-X_{(j-1)m+i}}{h_{n}}\right)1_{\left\{Y_{(j-1)m+i}>su_{n}^{x}\right\}},$
$R_{n}(s)$ $=\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}.$
@@
$\mathbb{E}\left[\widetilde{V}^{x,i}_{n}(s)\right]=\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}s\right\}}K\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\to s^{-1/\gamma(x)}g(x),$
@@
$\operatorname{var}\left(\widetilde{V}^{x,i}_{n}(s)\right)=\frac{\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor m}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\Bigg[\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}s\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}-\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}s\right\}}K\left(\frac{x-X}{h_{n}}\right)\right]^{2}}{h_{n}\overline{F}^{x}(u_{n}^{x})}\Bigg],$
@@
$\mathbb{E}\left[R_{n}(s)\right]=\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\frac{1}{n}\frac{\mathbb{E}\left[K\left(\frac{x-X}{h_{n}}\right)1_{\left\{Y>su_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\to 0,$
$\mathbb{E}\left[R_{n}(s)^{2}\right]=\frac{1}{\left(nh_{n}\overline{F}^{x}(u_{n}^{x})\right)^{2}}\biggl(\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]$
$\qquad\qquad\quad\ \ +2\sum_{j=\left\lfloor\frac{n}{m}\right\rfloor m}^{n}\sum_{i=j}^{n}\mathbb{E}\left[1_{\left\{Y_{i}>u_{n}^{x}s\right\}}K\left(\frac{x-X_{i}}{h_{n}}\right)1_{\left\{Y_{j}>u_{n}^{x}s\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)\right]\biggr)$
$\leq\frac{\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)}{\left(nh_{n}\overline{F}^{x}(u_{n}^{x})\right)^{2}}\biggl(\mathbb{E}\left[K^{2}\left(\frac{x-X}{h_{n}}\right)1_{\left\{Y>su_{n}^{x}\right\}}\right]+2\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\sqrt{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}s\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]^{2}}\ \biggr)$
$=\frac{\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)}{n^{2}h_{n}\overline{F}^{x}(u_{n}^{x})}\frac{\mathbb{E}\left[1_{\left\{Y>u_{n}^{x}t\right\}}K^{2}\left(\frac{x-X}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\left(1+2\left(n-\left\lfloor\frac{n}{m}\right\rfloor m\right)\right)\to 0,$
@@
[…]
@@
$\text{var}(\widetilde{V}_{n}^{x}(s))$ $=\frac{1}{(nh_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\Bigg(\sum_{j=1}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]-n^{2}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]^{2}$
$\qquad\qquad\qquad+2n\sum_{j=1}^{n}\left(1-\frac{j}{n}\right)\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]\Bigg)$
$\leq\frac{1}{(nh_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\Bigg(\sum_{j=1}^{n}\mathbb{E}\left[K^{2}\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]$
$\quad+2n\sum_{j=1}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right)\Bigg),$
@@
$\frac{1}{h_{n}\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]$
$=h_{n}\int_{[-1,1]^{2}}\biggl\{\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\mid X_{0}=x+w_{1}h_{n},X_{j}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr\}\mathrm{d}(w_{1},w_{2})=\mathcal{O}(h_{n}),$
where the integral is bounded, for $n$ large enough, due to condition 6 and smoothness of $g_{X_{0},X_{j}}$. Now, Billingsley’s inequality with the $\alpha$-mixing assumption yields that
$\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right)\right|$
$\leq 4\alpha(j)\left\lVert K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right\rVert_{\infty}\left\lVert K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right\rVert_{\infty}=\mathcal{O}(\alpha(j)),$
using that the kernel is bounded. Let $d_{n}=\lceil h_{n}^{-c_{x}}\rceil$ for some $c_{x}>0$. Combing the above considerations yields that
$\frac{2n}{(nh_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\left|\sum_{j=1}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}},K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right)\right|$
$\leq\frac{\text{cst}}{n(h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\left(\sum_{j=1}^{d_{n}-1}h_{n}^{2}\overline{F}^{x}(u_{n}^{x})+\sum_{j=d_{n}}^{\infty}j^{-\eta}\right)\leq\frac{\text{cst}}{n(h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\left(d_{n}h_{n}^{2}\overline{F}^{x}(u_{n}^{x})+d_{n}^{-\eta+1}\right)$
$=\text{cst}\left(\frac{d_{n}}{n\overline{F}^{x}(u_{n}^{x})}+\frac{d_{n}^{-\eta+1}}{n(h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}}\right)\sim\text{cst}\left(\frac{1}{nh_{n}^{c_{x}}\overline{F}^{x}(u_{n}^{x})}+\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}^{x}))^{2}}\right)\to 0.$